OpenStudy (anonymous):

look at these equations 48=3x2a 56=7x2a what are the values of a,b show your working out show sh

OpenStudy (bahrom7893):

48 = 3*2a 48 = 6a a = 48/6 = 8 56=7*2a 56 = 14a a = 56/14 a = 4 Please click on Good Answer if I helped. You can also hover over my name and click on Become a fan ;)!

OpenStudy (anonymous):

whats b

OpenStudy (bahrom7893):

u never gave an equation for b.

OpenStudy (anonymous):

I'm tearing my eyes for "b" in your question?, zaynelover

OpenStudy (bahrom7893):

maybe b = 4

OpenStudy (anonymous):

Does 3x2a mean 3 x^2?

OpenStudy (anonymous):

If there're no x variable, I'm suppose: 48=3x2a --> a = 48/ 6 = 8 56=7x2b ---> b = 56/14 = 4